OpenStudy (anonymous):

solve 2sec^2(x/2) - 3sec(x/2) - 2 =0

hero (hero):

So you have to solve 2sec^2(x/2) - 3sec(x/2) - 2 =0 for x, right?

OpenStudy (anonymous):

findthe general solution

hero (hero):

Well, for the moment, we can let y = sec^2(x/2) and if we did that, we would have 2y^2 - 3y - 2 = 0 Right?

OpenStudy (anonymous):

And then factor? (2y+1)(y-2)

hero (hero):

So after factoring we have (2y + 1)(y - 2) = 0 And from here, we can set 2y + 1 = 0 and y - 2 = 0 right?

OpenStudy (anonymous):

Alright I got that part now, so i continued solving and ended up getting sec(x/2) = -(.5) Which is no solution and and sec(x/2) = 2 which is where im now stuck

hero (hero):

So you're left with sec(x/2) = 2 From here, you can take the \(\sec^{-1}\) of both sides right?

OpenStudy (anonymous):

I don't have a calculator to do that, have to look off a unit circle

hero (hero):

How would you write the equation after taking the inverse secant of both sides?

OpenStudy (anonymous):

idk what you mean but im at cos(x/2) = pi/3 + 2pi and cos(x/2) = 5pi/3 +2pi after switching to cos and looking at the unit circle

hero (hero):

If you took the inverse secant of both sides, you would get x/2 = \(\sec^{-1}(2)\) right?

OpenStudy (anonymous):

yeah

hero (hero):

Are you able to find \(\sec^{-1}(2)\)?

OpenStudy (anonymous):

I think i got it thanks for taking me through the problem

hero (hero):

Just to be sure, what did you get for your final result?

hero (hero):

By the way, \(\sec^{-1}(x) \ne \cos(x)\)